# TBStar-Edit: From Image Editing Pattern Shifting to Consistency Enhancement

### v2

@@ 5 Conclusion
In this report, a new model named TBStar-Edit was proposed, focusing on high product-level consistency for e-commerce applications. Accurate and faithful image editing was achieved, resulting in significant improvements in model performance within high-consistency e-commerce scenarios. To address the issue of low-quality open-source data, a comprehensive data construction pipeline was introduced, encompassing data collection, construction, filtering, and augmentation. Through this pipeline, a large amount of high-quality data was generated, characterized by precise instruction following, strong object consistency, and superior image quality. Based on the concept of enhancing editing capability before reinforcing consistency, a hierarchical model framework was developed, consisting of a base model, pattern shifting modules, and consistency enhancement modules. A two-stage training strategy was adopted, enabling T2I models to acquire editing abilities and substantially improve consistency. Experimental results demonstrated that, in high-consistency e-commerce scenarios, the proposed model significantly outperformed current general SOTA models.

### v3

@@ 5 Conclusion
In this report, we introduce TBStar-Edit, a novel model designed to achieve high product-level consistency in e-commerce image editing. To overcome the limitations of low-quality open-source data, we develop a comprehensive data pipeline that yields a large-scale, high-quality dataset characterized by precise instruction adherence and strong subject preservation. Guided by the principle of "first establish editing capability then enhance consistency“, we design a hierarchical framework composed of a base model, pattern shifting modules, and consistency enhancement modules. This architecture is trained with a two-stage strategy that first equips the model with general editing abilities and then specifically refines its consistency. Experimental results demonstrate that TBStar-Edit significantly outperforms current state-of-the-art general-domain models, particularly in high-consistency e-commerce scenarios.
Appendix
6 More Visualization Results
Figure 6: More visualization results of TBStar-Edit.
Figure 7: More visualization results of TBStar-Edit.
Figure 8: More visualization results of TBStar-Edit.